# Special divisors in special divisor classes on graphs

### v2

@@ Abstract.
As in algebraic geometry, an effective divisor class on a vertex-weighted graph is called special if also its residual class is effective. We study the question, when this is true already on the level of divisors; that is, when there exists an effective divisor in the class whose residual is effective as well, so called uniform divisors. We show that uniform divisors exist in any special class on graphs with all vertex weights non-zero; we show that in general a representative can be chosen, whose value is off by at most one on vertices of weight zero. To prove this, we generalize the notion of divisors that are reduced with respect to a vertex to divisors that are reduced with respect to a set of vertices. We compare this notion to a previous generalization due to Luo. As an application, we obtain a new algorithm to determine whether a divisor class is effective.
The author was partially supported by NSF FRG grant DMS–2053261.
@@ 1. Introduction
The study of divisors on algebraic curves is one of the classic topics of algebraic geometry. More recently, a divisor theory for finite graphs has been developed that shares striking similarities with the theory for algebraic curves. Starting with [BN07], these analogies have been studied extensively; see, for example, [MZ08], [Bak08], [CV10], [AC13], [CLM15], [vDdBG20], [MUW21], [CPS23] and [CM23].
Concretely, a divisor $\underline{d}$ on a vertex-weighted graph $G$ is a linear combination of vertices of $G$. The basic connection to algebraic geometry is given by viewing the multidegree of a line bundle on a nodal curve with dual graph $G$ as a divisor on $G$.
As in algebraic geometry, one defines an equivalence relation on divisors, called linear equivalence, and there is a canonical divisor class $\omega$ on each graph $G$. The divisor classes are given by divisors that differ from each other by a series of chip-firing moves. A divisor class $\delta$ is called special if both $\delta$ and its residual class $\delta^{*}\coloneqq\omega-\delta$ are effective. That is, both $\delta$ and $\delta^{*}$ need to contain a divisor that has positive value on each vertex. See Section 2 for a more detailed exposition of these notions.
Different from algebraic geometry, the canonical class $\omega$ on $G$ contains a distinguished representative, which allows the definition of a residual already on the level of divisors. It is the divisor $\underline{k}\in\omega$ with value
$\underline{k}_{v}=2g_{v}-2+\operatorname{val}(v)$
on a vertex $v$ of $G$; here $g_{v}$ is the weight of the vertex and $\operatorname{val}(v)$ its valence. The divisor $\underline{k}$ is the multidegree of the dualizing sheaf on a nodal curve with dual graph $G$. In this way, one can canonically define the residual of a divisor $\underline{d}$ on $G$ as $\underline{d}^{*}\coloneqq\underline{k}-\underline{d}$.
@@
If both $\underline{d}$ and its residual $\underline{d}^{*}$ are effective, we call $\underline{d}$ a uniform divisor. Concretely, this means a divisor $\underline{d}$ whose value $\underline{d}_{v}$ on any vertex $v\in G$ satisfies
@@
In algebro-geometric terms, this translates to a multidegree that lies between that of the structure sheaf and the dualizing sheaf.
The class of the residual divisor $\underline{d}^{*}$ is the residual class $\delta^{*}$. Thus if $\underline{d}$ is uniform, clearly the class $\delta$ of $\underline{d}$ is special. The converse does not hold, and there are special classes which contain no uniform representatives (see Example 5.3). We show however, that in any special class there are representatives that are almost uniform in the following sense (see Theorem 5.5):
Theorem A.
@@
In particular, if every vertex of $G$ with weight 00 is adjacent to a loop, then every special divisor class contains a uniform representative.
While the theorem is purely combinatorial, our original motivation for studying this question is algebro-geometric: Together with results in [Chr23], Theorem A allows to construct explicit Clifford representatives for a large class of graphs $G$; that is, divisors $\underline{d}\in\delta$, such that every line bundle of multidegree $\underline{d}$ satisfies the classical Clifford inequality. We explore this and related questions in [BCM24].
To prove Theorem A, we study $V$-reduced divisors for a subset of vertices $V$, which were introduced independently by Barbosa [Bar21]. They generalize the $v$-reduced divisors for a single vertex $v$, which are one of the most important tools in the study of divisors on graphs; see, for example, [BN07], [Luo11], [CDPR12], [Ami13], [HKN13], [CLM15].
Definition 1.1.
A divisor $\underline{d}$ is $V$-reduced if it is effective away from $V$ and a chip-firing move along any set of vertices $W$ with $W\cap V=0$ turns $\underline{d}$ negative on at least one vertex in $W$.
Different from $v$-reduced divisors, $V$-reduced divisors need not be unique in their divisor class. But $V$-reduced divisors do share many important properties with the $v$-reduced ones and we use them to give a new algorithm to produce an effective divisor linearly equivalent to a given divisor $\underline{d}$, or to certify that no such effective divisor exists. It is a modification of the commonly used algorithm due to Baker and Shokrieh [BS13, Algorithm 4], which skips the step in which a divisor is constructed that is effective away from $v$ and instead directly proceeds by calculating successively $V_{i}$-reduced representatives for a decreasing chain of sets of vertices $V_{i}$ (see Algorithm 3.10). We use this algorithm in turn to produce the representatives in Theorem A.
In a different direction, we compare the notion of $V$-reduced divisors to a different generalization of $v$-reduced divisors introduced by Luo [Luo13] (see Theorem 4.9):
Theorem B.
1.1. Structure of the paper
In Section 2, we recall basic definitions concerning graphs and divisors. In Section 3, we introduce the Dhar decomposition with respect to a subset of vertices $V\subset V(G)$ and $V$-reduced divisors. The section also contains Algorithm 3.10, which is central for the proof of Theorem A. In Section 4, we discuss the relation of $V$-reduced divisors to the $\mathcal{E}$-reduced divisors of Luo proving Theorem B. In Section 5, we study the existence of uniform representatives in special divisor classes to prove Theorem A.
Acknowledgements. I would like to thank Sam Payne for pointing out Luo’s work on $\mathcal{E}$-reduced divisors, and Matt Baker for comments on an earlier version of the paper. The notion of $V$-reduced divisors was introduced independently by Myrla Barbosa in her PhD thesis [Bar21], which was pointed out by Margarida Melo. I’m grateful to both of them for helpful discussions related to this circle of ideas.
@@ 2.3. Linear equivalence
where the sum runs over all edges adjacent to $v$. An $A$-divisor of the form $\mathrm{div}(f)$ has always degree 00 and is called a principal divisor. Two functions $f,g\in C^{0}(G;A)$ have the same associated $A$-divisor if and only if one function is obtained from the other by adding the same constant at each vertex.
@@
The set of equivalence classes under this relation is denoted by ${\operatorname{{Pic}}}(G)$. Since principal divisors have degree 00, the classes $\delta\in{\operatorname{{Pic}}}(G)$ have a well-defined degree given by the degree of its representatives. We call a class $\delta\in{\operatorname{{Pic}}}(G)$ effective, if $\delta$ contains an effective representative.
For a subset of vertices $V\subset V(G)$ we denote by $\mathbb{1}_{V}$ the characteristic function of $V$; that is, $\mathbb{1}_{V}$ has value $1$ on all vertices in $V$ and value 00 on all other vertices. Given a divisor $\underline{d}\in{\operatorname{{Div}}}(G)$, we say that the linearly equivalent divisor
@@ 3.2. $V$-reduced divisors
Changing $f$ by adding to its value the same integer at each vertex of $G$ does not change $\mathrm{div}(f)$. Thus we may assume that the minimal value of $f$ on vertices of $G$ is 00. Since $\underline{d}$ itself is not effective, $f$ is not zero on all vertices of $G$.
@@
where $\mathbb{1}_{F_{i}}$ has value $1$ on vertices in $F_{i}$ and 00 otherwise. Since $\mathrm{div}(\mathbb{1}_{F_{i}})$ encodes a chip-firing move along $F_{i}$, this means that $\underline{d}^{\prime}$ is obtained from $\underline{d}$ by successively chip-firing along the $F_{i}$.
@@ 4.1. $\mathcal{E}$-reduced divisors
and the minimal value of $q_{\mathcal{E}}(\mathcal{D})$ equals 00.
@@ 4.2. Preliminary observations
Before we can clarify the relation between $\mathcal{E}$- and $V$-reduced divisors, we need to collect some auxiliary results. Recall that for $f\in C^{0}(G;\mathbb{R})$ we denote by $U_{0}(f)$ the set of vertices on which $f$ has value 00.
@@
and the minimal value of $f$ is 00. Assume that $U_{0}(f)\cap U_{0}\left(q_{\mathcal{E}}(\underline{d}^{\prime})\right)\neq\emptyset$. Then
@@
Since $U_{0}(f)\cap U_{0}(q_{\mathcal{E}}(\underline{d}^{\prime}))\neq\emptyset$, the minimal value of $q_{\mathcal{E}}(\underline{d}^{\prime})+f$ is 00. Thus by definition
@@
Denote by $\mathbb{1}_{V^{\prime}}$ the characteristic function of $V^{\prime}$ (i.e. with value $1$ on each vertex in $V^{\prime}$ and value 00 on all other vertices). By definition, we then have
@@
    and for all vertices $u\neq v$ and $u\not\in V$, the sum of the outgoing slopes of $f$ at $u$ is 00, i.e.,
@@
For any $v_{i}\in V$ let $q_{v}(v_{i})\in C(G;\mathbb{R})$ denote the unique function such that $v=v_{i}+{\operatorname{{div}}}\left(q_{v}(v_{i})\right)$ and such that the minimal value of $q_{v}(v_{i})$ is 00, which is achieved at $v$. By construction, for any vertex $u$ different from $v$ and $v_{i}$, the sum of the outgoing slopes of $q_{v}(v_{i})$ at $u$ is zero. Thus any linear combination
@@
The third claim in Lemma 4.5 in fact does not just hold for vertices $u\neq v$ and $u\not\in V$, but holds also for vertices $u\in V$ such that there is no simple path from $v$ to $u$ that does not contain any other vertex of $V$. Indeed, the first and third claim imply that $f$ needs to have constant value $k$ on any connected component of $G-V$ that does not contain $v$. Hence $f$ has value $k$ on all vertices adjacent to a vertex $u$ as above. Since $f$ has value $k$ also on $u$, this implies the third claim.
@@
    If $v\in U_{0}(q_{\mathcal{E}}(\mathcal{D}))$ and $v\not\in V$, then $U_{0}(q_{\mathcal{E}}(\mathcal{D}))$ contains the entire connected component $G^{\prime}$ of $G-V$ that contains $v$, as well as any vertex in $V$ adjacent to a vertex in $G^{\prime}$.
@@
Recall that $q_{\mathcal{E}}(\mathcal{D})$ is normalized so that the vertices at which it is 00 are precisely those at which it achieves its minimum value and that we have by definition
@@
Thus the sum of the outgoing slopes of $q_{\mathcal{E}}(\mathcal{D})$ at $v$ is 00. Since $v$ is assumed to be a local minimum of $q_{\mathcal{E}}(\mathcal{D})$, this shows that the value of $q_{\mathcal{E}}(\mathcal{D})$ needs to be the same on $v$ as on all vertices adjacent to $v$. Repeating this procedure for vertices in $V(G)\setminus V$ that are adjacent to $v$ shows the second claim. For the first claim, observe that since $G$ is connected, eventually a vertex contained in $V$ needs to be adjacent to the vertex $v$. ∎
@@
Let $V\subset V(G)$ be a set of vertices and $\mathcal{D}$ an $\mathbb{R}$-divisor of degree $d$, effective away from $V$. Then there exists an $\mathbb{R}$-divisor $\mathcal{E}$ of degree $d$, supported on $V$ and such that $q_{\mathcal{E}}(\mathcal{D})$ has value 00 on all vertices in $V$.
@@
For the inductive step, let $v\in V-U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$ be a vertex in $V$ on which $q_{\mathcal{E}}(\mathcal{D})$ is not zero and hence positive. By Lemma 4.5 there exists an $\mathbb{R}$-valued function $f_{k}\in C(G;\mathbb{R})$ that has value $k$ on $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$, value 00 on $v$ and such that the sum of the outgoing slopes of $f_{k}$ is 00 at any vertex $u$ different from $v$ and not contained in $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$. Set
@@
Since $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$ is the set of vertices on which $q_{\mathcal{E}}(\mathcal{D})$ has minimal value (equal to 00), $q_{\mathcal{E}}+f_{k}$ still has minimal value (equal to $k$) on $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$ as long as $k$ is positive but small enough. It follows that

### v3

@@ Abstract.
We generalize the notion of divisors that are reduced with respect to a vertex to divisors that are reduced with respect to a set of vertices. We establish properties that remain valid in this more general context which allow us to improve the standard algorithm to determine whether a divisor class is effective. We then characterize reduced divisors in our sense in terms of Luo’s potential-theoretic generalization of reduced divisors. Finally, as a further application, we use this setup to study the existence of so called uniform representatives in special divisor classes.
The author was partially supported by NSF FRG grant DMS–2053261 and is member of the INdaM group GNSAGA.
@@ 1. Introduction
The study of divisors on algebraic curves is one of the classic topics of algebraic geometry. More recently, a divisor theory for finite graphs has been developed that shares striking similarities with the theory for algebraic curves. Starting with [BN07], these analogies have been studied extensively and partially explained; see, for example, [MZ08, Bak08, CV10, AC13, CLM15, AP20, FJP20, MUW21, CPS23].
Concretely, a divisor $\underline{d}$ on a vertex-weighted graph $G$ is simply a $\mathbb{Z}$-linear combination of vertices of $G$. Two divisors on $G$ are called linearly equivalent if one can be obtained from the other by a series of so-called chip-firing moves. Thinking about $\underline{d}$ as a collection of chips on each vertex, a chip-firing move based at a vertex $v$ is given by moving one chip of $\underline{d}$ along each edge adjacent to $v$.
For a vertex $v$ of $G$ there exists a unique $v$-reduced divisor in any linear equivalence class. Roughly speaking, it is a divisor with the minimal possible non-negative number of chips away from $v$. These representatives have been introduced already in [BN07] and remain one of the main tools in studying the Baker-Norine rank of divisors on graphs. One of the reasons for their usefulness in this context is that a divisor class contains an effective representative, i.e. one with all values non-negative, if and only if the $v$-reduced representative is effective.
In the current paper, we generalize the notion of $v$-reduced divisors by replacing the single vertex $v$ with a subset of vertices $V$. We show that some of the properties that make $v$-reduced divisors so useful generalize to this more general setting, and use them to improve the standard algorithm due to Baker and Shokrieh [BS13] to find effective representatives in chip-firing classes or certify that no such representatives exist. We then give a precise relation of our notion with a different generalization of $v$-reduced divisors due to Luo [Luo13]. Finally, we use $V$-reduced representatives to study the existence of so called uniform representatives in special divisor classes, that are used in the subsequent paper [BCM25] to give algebro-geometric applications.
Main results
Our main object of interest in this paper is the following generalization of $v$-reduced divisors (introduced independently by Myrla Barbosa [Bar21] in her PhD thesis):
Definition 1.1.
Let $V$ be a set of vertices of a graph $G$. A divisor $\underline{d}$ is $V$-reduced if it is effective away from $V$ and a chip-firing move along any set of vertices $W$ with $W\cap V=\emptyset$ turns $\underline{d}$ negative on at least one vertex in $W$.
Different from $v$-reduced divisors, $V$-reduced divisors need not be unique in their divisor class, and it is also not true that a divisor class contains an effective representative if and only if the $V$-reduced representatives are effective. But it is easy to see that every divisor class contains a $V$-reduced representative for any $V$ (Lemma 3.4). And the following proposition gives a relation with effective representatives, see Proposition 3.8:
Proposition 1.2.
$W_{\mathrm{Dhar}}(\underline{d},V)\neq\emptyset.$
Here $W_{\mathrm{Dhar}}(\underline{d},V)$ denotes the Dhar decomposition of $G$ with respect to $\underline{d}$ and $V$ which we introduce in Section 3.1. Roughly speaking, it is the maximal set of vertices along which a chip-firing move of $\underline{d}$ does not create any new negative values for $\underline{d}$.
We use Proposition 1.2 to give a new algorithm to produce an effective divisor linearly equivalent to a given divisor $\underline{d}$, or to certify that no such effective divisor exists. It is a modification of the commonly used algorithm due to Baker and Shokrieh [BS13, Algorithm 4], which skips the step in which a divisor is constructed that is effective away from $v$ and instead directly proceeds by calculating successively $V_{i}$-reduced representatives for a decreasing chain of sets of vertices $V_{i}$ obtained by chip-firing along the $W_{\mathrm{Dhar}}(\underline{d},V_{i})$; see Algorithm 3.10 for details.
Next, we compare the notion of $V$-reduced divisors to a different generalization of $v$-reduced divisors introduced by Luo [Luo13] that generalizes the potential-theoretic characterization of $v$-reduced divisors given in [BS13], see Theorem 4.9:
Theorem A.
Here an $\mathbb{R}$-divisor is simply a divisor with values in $\mathbb{R}$ instead of $\mathbb{Z}$. An $\mathcal{E}$-reduced divisor on the other hand is a divisor which, roughly speaking, has minimal distance to $\mathcal{E}$ in its linear equivalence class; see Section 4 for details.
Finally, as an application, we study the existence of so called uniform divisors in linear equivalence classes of effective divisors whose residual is also effective. To this end, we need to introduce some notation. Recall that the canonical divisor $\underline{k}\in\omega$ is given by
$\underline{k}_{v}=2g_{v}-2+\operatorname{val}(v),$
where $g_{v}$ is the weight of the vertex $v$ and $\operatorname{val}(v)$ its valence. We denote the linear equivalence class of $\underline{k}$ by $\omega$.
@@
Given a divisor $\underline{d}$ in the equivalence class $\delta$, its residual divisor is defined as $\underline{d}^{*}\coloneqq\underline{k}-\underline{d}$. Similarly, its residual class is $\delta^{*}\coloneqq\omega-\delta$. A divisor class $\delta$ is called special if both $\delta$ and its residual class $\delta^{*}\coloneqq\omega-\delta$ are effective. That is, both $\delta$ and $\delta^{*}$ need to contain a divisor that has non-negative value on each vertex. If both $\underline{d}$ and its residual $\underline{d}^{*}$ are effective, we call $\underline{d}$ a uniform divisor. Concretely, this means a divisor $\underline{d}$ whose value $\underline{d}_{v}$ on any vertex $v\in G$ satisfies
@@
If $\underline{d}$ is uniform, clearly the class $\delta$ of $\underline{d}$ is special. The converse does not hold, and there are special classes which contain no uniform representatives (see Example 5.3). We show however, that in any special class there are representatives that are almost uniform in the following sense (see Theorem 5.6):
Theorem B.
@@
In particular, if every vertex of $G$ with weight 0 is adjacent to a loop, then every special divisor class contains a uniform representative.
While the theorem is purely combinatorial, our original motivation for studying this question is algebro-geometric: Together with results of [Chr23], Theorem A allows to construct explicit Clifford representatives for a large class of graphs $G$; that is, divisors $\underline{d}\in\delta$, such that every line bundle of multidegree $\underline{d}$ satisfies the classical Clifford inequality. We explore this and related questions in [BCM25].
Acknowledgements. I would like to thank Sam Payne for pointing out Luo’s work on $\mathcal{E}$-reduced divisors, and Matt Baker for comments on an earlier version of the paper. The notion of $V$-reduced divisors was introduced independently by Myrla Barbosa in her PhD thesis [Bar21], which was pointed out by Margarida Melo. I’m grateful to both of them for helpful discussions related to this circle of ideas. Finally, I would like to thank an anonymous referee for pointing out the alternative proof of Theorem B using partial orientations explained in Remark 5.8.
@@ 2.3. Linear equivalence
where the sum runs over all edges adjacent to $v$. An $A$-divisor of the form $\mathrm{div}(f)$ has always degree 0 and is called a principal divisor. Two functions $f,g\in C^{0}(G;A)$ have the same associated $A$-divisor if and only if one function is obtained from the other by adding the same constant at each vertex.
@@
The set of equivalence classes under this relation is denoted by ${\operatorname{{Pic}}}(G)$. Since principal divisors have degree 0, the classes $\delta\in{\operatorname{{Pic}}}(G)$ have a well-defined degree given by the degree of its representatives. We call a class $\delta\in{\operatorname{{Pic}}}(G)$ effective, if $\delta$ contains an effective representative.
For a subset of vertices $V\subset V(G)$ we denote by $\mathbb{1}_{V}$ the characteristic function of $V$; that is, $\mathbb{1}_{V}$ has value $1$ on all vertices in $V$ and value 0 on all other vertices. Given a divisor $\underline{d}\in{\operatorname{{Div}}}(G)$, we say that the linearly equivalent divisor
@@ 3.2. $V$-reduced divisors
Changing $f$ by adding to its value the same integer at each vertex of $G$ does not change $\mathrm{div}(f)$. Thus we may assume that the minimal value of $f$ on vertices of $G$ is 0. Since $\underline{d}$ itself is not effective, $f$ is not zero on all vertices of $G$.
@@
where $\mathbb{1}_{F_{i}}$ has value $1$ on vertices in $F_{i}$ and 0 otherwise. Since $\mathrm{div}(\mathbb{1}_{F_{i}})$ encodes a chip-firing move along $F_{i}$, this means that $\underline{d}^{\prime}$ is obtained from $\underline{d}$ by successively chip-firing along the $F_{i}$.
@@ 4.1. $\mathcal{E}$-reduced divisors
and the minimal value of $q_{\mathcal{E}}(\mathcal{D})$ equals 0.
@@ 4.2. Preliminary observations
Before we can clarify the relation between $\mathcal{E}$- and $V$-reduced divisors, we need to collect some auxiliary results. Recall that for $f\in C^{0}(G;\mathbb{R})$ we denote by $U_{0}(f)$ the set of vertices on which $f$ has value 0.
@@
and the minimal value of $f$ is 0. Assume that $U_{0}(f)\cap U_{0}\left(q_{\mathcal{E}}(\underline{d}^{\prime})\right)\neq\emptyset$. Then
@@
Since $U_{0}(f)\cap U_{0}(q_{\mathcal{E}}(\underline{d}^{\prime}))\neq\emptyset$, the minimal value of $q_{\mathcal{E}}(\underline{d}^{\prime})+f$ is 0. Thus by definition
@@
Denote by $\mathbb{1}_{V^{\prime}}$ the characteristic function of $V^{\prime}$ (i.e. with value $1$ on each vertex in $V^{\prime}$ and value 0 on all other vertices). By definition, we then have
@@
    and for all vertices $u\neq v$ and $u\not\in V$, the sum of the outgoing slopes of $f$ at $u$ is 0, i.e.,
@@
For any $v_{i}\in V$ let $q_{v}(v_{i})\in C(G;\mathbb{R})$ denote the unique function such that $v=v_{i}+{\operatorname{{div}}}\left(q_{v}(v_{i})\right)$ and such that the minimal value of $q_{v}(v_{i})$ is 0, which is achieved at $v$. By construction, for any vertex $u$ different from $v$ and $v_{i}$, the sum of the outgoing slopes of $q_{v}(v_{i})$ at $u$ is zero. Thus any linear combination
@@
The third claim in Lemma 4.5 in fact does not just hold for vertices $u\neq v$ and $u\not\in V$, but holds also for vertices $u\in V$ such that there is no simple path from $v$ to $u$ that does not contain any other vertex of $V$. Indeed, the first and third claim imply that $f$ needs to have constant value $k$ on any connected component of GVG-V that does not contain $v$. Hence $f$ has value $k$ on all vertices adjacent to a vertex $u$ as above. Since $f$ has value $k$ also on $u$, this implies the third claim.
@@
    If $v\in U_{0}(q_{\mathcal{E}}(\mathcal{D}))$ and $v\not\in V$, then $U_{0}(q_{\mathcal{E}}(\mathcal{D}))$ contains the entire connected component $G^{\prime}$ of GVG-V that contains $v$, as well as any vertex in $V$ adjacent to a vertex in $G^{\prime}$.
@@
Recall that $q_{\mathcal{E}}(\mathcal{D})$ is normalized so that the vertices at which it is 0 are precisely those at which it achieves its minimum value and that we have by definition
@@
Thus the sum of the outgoing slopes of $q_{\mathcal{E}}(\mathcal{D})$ at $v$ is 0. Since $v$ is assumed to be a local minimum of $q_{\mathcal{E}}(\mathcal{D})$, this shows that the value of $q_{\mathcal{E}}(\mathcal{D})$ needs to be the same on $v$ as on all vertices adjacent to $v$. Repeating this procedure for vertices in $V(G)\setminus V$ that are adjacent to $v$ shows the second claim. For the first claim, observe that since $G$ is connected, eventually a vertex contained in $V$ needs to be adjacent to the vertex $v$. ∎
@@
Let $V\subset V(G)$ be a set of vertices and $\mathcal{D}$ an $\mathbb{R}$-divisor of degree $d$, effective away from $V$. Then there exists an $\mathbb{R}$-divisor $\mathcal{E}$ of degree $d$, supported on $V$ and such that $q_{\mathcal{E}}(\mathcal{D})$ has value 0 on all vertices in $V$.
@@
For the inductive step, let $v\in V-U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$ be a vertex in $V$ on which $q_{\mathcal{E}}(\mathcal{D})$ is not zero and hence positive. By Lemma 4.5 there exists an $\mathbb{R}$-valued function $f_{k}\in C(G;\mathbb{R})$ that has value $k$ on $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$, value 0 on $v$ and such that the sum of the outgoing slopes of $f_{k}$ is 0 at any vertex $u$ different from $v$ and not contained in $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$. Set
@@
Since $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$ is the set of vertices on which $q_{\mathcal{E}}(\mathcal{D})$ has minimal value (equal to 0), $q_{\mathcal{E}}+f_{k}$ still has minimal value (equal to $k$) on $U_{0}\left(q_{\mathcal{E}}(\mathcal{D})\right)$ as long as $k$ is positive but small enough. It follows that